# Do Automatic Factuality Metrics Measure Factuality? A Critical Evaluation

### v3

@@ Abstract
Modern LLMs can now produce highly readable abstractive summaries, to the point where traditional automated metrics for evaluating summary quality, such as ROUGE, have become saturated. However, LLMs still sometimes introduce unwanted content into summaries, i.e., information inconsistent with or unsupported by their source. Measuring the occurrence of these often subtle “hallucinations” automatically has proved to be challenging. This in turn has motivated development of a variety of metrics intended to measure the factual consistency of generated summaries against their source. But are these approaches measuring what they purport to do? In this work, we stress-test automatic factuality metrics. Specifically, we investigate whether and to what degree superficial attributes of summary texts suffice to predict “factuality”, finding that a (supervised) model using only such shallow features is reasonably competitive with SOTA factuality scoring methods. We then evaluate how factuality metrics respond to factual corrections in inconsistent summaries and find that only a few show meaningful improvements. In contrast, some metrics are more sensitive to benign, non-factual edits. Motivated by these insights, we show that one can “game” (most) automatic factuality metrics, i.e., reliably inflate “factuality” scores by appending innocuous sentences to generated summaries. Taken together, our results raise questions about the degree to which we should rely on existing automated factuality metrics and what exactly we want “factuality metrics” to measure.
Do Automatic Factuality Metrics Measure Factuality?
A Critical Evaluation
Sanjana Ramprasad Northeastern University ramprasad.sa@northeastern.edu                        Byron C. Wallace Northeastern University b.wallace@northeastern.edu
Figure 0.1: Many methods have been proposed to automatically evaluate the faithfulness (or factual consistency) of summaries with respect to inputs. In this work we critically evaluate such approaches, e.g., by measuring their sensitivity to various manipulations, as shown here.
@@ 1 Introduction
LLMs are generally strong abstractive summarizers Goyal et al. (2022); Zhang et al. (2024). Still, they are not infallible. Even the largest, most capable models sometimes introduce subtle “hallucinations” (or “confabulations”) into summaries that are unsupported by or in contradiction to the corresponding input document (Zhang et al., 2024; Tang et al., 2024b; Ramprasad et al., 2024b). Such behaviour is especially problematic in domains such as medicine or law, where inaccurate information could translate into meaningfully negative consequences for individuals.
@@
However, manually evaluating model outputs’ faithfulness with respect to their source is expensive, time-consuming, and impractical to scale. This has motivated the development of automated methods that score generated summaries for faithfulness with respect to reference documents. Such metrics have been operationalized using a range of underlying techniques, including entailment (Laban et al., 2022; Scirè et al., 2024), QA models (Scialom et al., 2021; Fabbri et al., 2021b), specialized models explicitly trained to score source-summary pairs (Zhong et al., 2022; Zha et al., 2023; Tang et al., 2024a), and more recently, prompt-based methods that rely on LLM calls (Luo et al., 2023; Wang et al., 2023a).
@@
Most of these metrics have been evaluated against human benchmark assessments (binary labels or Likert scores) of factual consistency (Maynez et al., 2020; Fabbri et al., 2021a; Laban et al., 2022; Honovich et al., 2022; Wang et al., 2022; Gao et al., 2023; Tang et al., 2024a). These assessments have established that automated factuality metrics correlate with human evaluations to varying degrees. But are such approaches actually attuned to subtle question of factual consistency between inputs and outputs, or are they merely relying on shallow heuristic signals—such as lexical overlap, entity repetition, or the presence of specific phrases—to make their judgments? For instance, Kamoi et al. (2023b) assessed the reliability of QA-based metrics and found that while they can predict summary-level factuality, they fail to identify correct error spans, and are outperformed by simple exact match baselines.
If metrics do rely on shallow heuristic patterns, this raises serious concerns about the reliability of these metrics, making them vulnerable to “gaming”, where summaries or claims can be crafted to exploit heuristic shortcuts and artificially inflate scores without genuinely improving factual accuracy.
In this work, we stress test a diverse set of SOTA factual consistency metrics on context-summary pairs that yield continuous factual consistency scores. Our analyses offer the following empirical results as our main contributions.
Superficial attributes of text predict “factuality” nearly as well as complex SOTA metrics when supervised. We define a small set of straightforward heuristic predictors (e.g. lexical overlap, conciseness, word novelty etc.) By themselves, such features do not directly imply factual consistency (or a lack thereof). But we find that fine-tuning a simple MLP defined over these features typically yields performance competitive with (often complex) SOTA factual consistency metrics. This suggests that it is at least possible that existing metrics rely (merely) on superficial cues which correlate with consistency, rather than making a nuanced assessment of accuracy.
Factuality metrics are somewhat responsive to factual corrections, but also often sensitive to irrelevant (benign) modifications to summaries. We use a previously released dataset (Krishna et al., 2024) comprising a small set of summaries labelled as being inconsistent with the corresponding input, paired with minimally edited versions that fix these inconsistencies. We use this set to assess whether factuality metrics are responsive to these edits, i.e., whether the scores they assign to the faithful versions of summaries are higher than the original inconsistent versions. We find that most metrics are at least somewhat responsive in this way. However, we also consider the complement of this experiment: We modify summaries in ways that should not affect factual consistency (e.g., paraphrasing texts) and measure changes in factuality scores assigned. Metrics would ideally be invariant to such benign perturbations, but we find that they are, on average, sensitive to them. In fact, some metrics exhibit greater responsiveness to benign changes in inaccurate summaries than to actual factual corrections, raising concerns about their reliability in distinguishing genuine improvements from inconsequential edits.
@@
(Some) factual consistency metrics are gameable. The preceding findings suggest that factuality metrics rely at least partially on superficial cues that do not directly reflect factual consistency. In turn, this suggests that we should be able to “game” them by inserting cues to systematically inflate scores assigned to model outputs. Indeed, we find that superfluous innocuous phrases used standalone or with summaries induce higher factuality scores. For many fine-tuned metrics, these artificial boosts from “gamed” summaries exceed the gains achieved from summaries generated by more factual models, exposing a critical vulnerability.
While previous efforts have meta-evaluated factuality metrics (Gabriel et al., 2021; Chen et al., 2021a), these approaches primarily focus on inducing errors to assess metrics’ sensitivity to specific error types or their general detection capabilities. In contrast, our work examines the degree to which factuality metrics respond to factual changes compared to unrelated textual attributes or artifacts, shedding light on their susceptibility to spurious correlations and non-factual influences.
@@ 2.1 Benchmark Datasets
Automatic factuality metrics for summarization have been evaluated using various benchmarks, predominantly based on news sources (Hermann et al., 2015; Narayan et al., 2018). The AggreFact dataset (Tang et al., 2022) consolidates several benchmarks on fine-tuned model-generated summaries from such sources (Maynez et al., 2020; Wang et al., 2020; Pagnoni et al., 2021; Fabbri et al., 2021a; Honovich et al., 2022; Laban et al., 2022). Similarly, datasets like FacEval (Wang et al., 2022) and ReferenceMatters (Gao et al., 2023) benchmark dialogue summarization of fine-tuned models on sources from SAMSum (Gliwa et al., 2019a) and DialogSum (Chen et al., 2021b).
However, these datasets primarily focus on summaries generated by fine-tuned models that predate modern LLMs, which have demonstrated superior zero-shot summarization capabilities (Goyal et al., 2022). As a result, the performance of factuality metrics on these benchmarks may not accurately reflect their effectiveness on outputs from SOTA LLMs, which are prone to different types of errors (Tang et al., 2022). To bridge this gap, Tang et al. (2024a) introduced the LLM-AggreFact dataset, which includes factuality labels for summaries generated by recent LLMs across a variety of domains, including Wikipedia, interviews, news, dialogues, science, and healthcare. Similarly, Krishna et al. (2024) released the GenAudit dataset with factuality annotations for LLM summaries in news, Reddit, and clinical settings, while Ramprasad et al. (2024a) focus on LLM generated dialogue summaries in the LLM-dialogue dataset.
For our analysis we use all of the above benchmarks to capture a wide range of error types. For fine-tuned model summaries, we use AggreFact for news and FacEval for dialogues. For LLM-generated summaries, we rely on LLM-AggreFact, GenAudit, and LLM-dialogue across varied domains. We note that each benchmark consolidates multiple datasets and to ensure clean separation of distributions, we avoid overlapping datasets between our test and development splits. We achieve this by randomly selecting datasets from each benchmark for the development set (we include only the dev split), while the remaining datasets are reserved for testing (using their respective test split). Given its smaller size, the GenAudit dataset is included exclusively in the test set. A detailed breakdown of the dev and test splits is provided in Appendix A.1.
@@ 2.2 Automatic Factuality Metrics for Summarization
For QA-based metrics we use QuestEval (Scialom et al., 2021). As an NLI-based metric we use SummaC-Conv (Laban et al., 2022). We use three specialized models for evaluation: UniEval, AlignScore and MiniCheck. UniEval Zhong et al. (2022) reframes NLG evaluation as a Boolean QA task and uses T5 (Raffel et al., 2020) to score different dimensions. AlignScore (Zha et al., 2023) evaluates summaries by combining an alignment function—a RoBERTa model (Liu, 2019) fine-tuned on diverse tasks—with a splitting and aggregation strategy. MiniCheck(Tang et al., 2024a) uses a Flan-T5 model (Chung et al., 2022) fine-tuned on a synthetic dataset created by the authors.
We also use GPT-4 to score the factual consistency of summaries based on a direct assessment (DA) prompt template from Wang et al. (2023a).
@@
QuestEval (Scialom et al., 2021) QA
SummaC-Conv (Laban et al., 2022) NLI
UniEval (Zhong et al., 2022) specialized model
AlignScore (Zha et al., 2023) specialized model
MiniCheck (Tang et al., 2024a) specialized model
ChatGPT-DA (Wang et al., 2023a) Prompt / LLM
@@
2.3 Heuristic Indicators of “Factuality”
Figure 2.1: We train a simple MLP (bottom bar) over only superficial indicators of consistency (e.g., word overlap). This realizes performance comparable to SOTA consistency metrics for in most cases. This model is supervised, but we train and test on disjoint datasets. Our results raise questions about how much inference is necessary to assess consistency on these benchmark corpora, given what superficial predictors reveal.
Another predictor we use is semantic similarity. Specifically, embed summary and source sentences via BERT (Devlin, 2018), then score each summary sentence based on which source sentence it is most similar to (using cosine similarity). We average all summary sentence scores to derive a final score.
3 Are Superficial Features Sufficient to Infer “Factuality”?
The degree to which a summary is faithful to the text it summarizes is a subtle question that demands understanding the content in both texts to determine if they are consistent. We would therefore expect that metrics capable of measuring factual consistency in general would need to capitalize on information beyond the superficial features of source and summary texts just discussed.
To assess if this is in fact the case we first evaluate the extent to which a model using superficial features alone can be trained to infer factuality. We train an MLP to predict (human) factuality labels as a function of shallow input features. We compare the results achieved—specifically, AUC scores—using this method to a suite of more sophisticated approaches that have been previously proposed.
The shallow model has been explicitly supervised. To replicate the scenario of metrics trained on different datasets, we ensure that our training set is entirely separate from the benchmark datasets in the test set; this makes the approach akin to the supervised, “specialized” models for measuring factual consistency considered.
We emphasize that the purpose of this exercise is not to propose a practical “new” approach to score summaries for factuality. Rather, our point here is to determine the extent to which existing metrics improve performance beyond what can be achieved using simple, superficial attributes of text.
Figure 2.1 summarizes our findings. Perhaps surprisingly, this shallow model is generally competitive with some previously proposed (complex) models. On document and query-focused summarization (QFS), the shallow model performs comparably to established QA/NLI metrics and the custom metric UniEval. In fact, it outperforms these metrics in the context of dialogue summarization evaluation. However, the shallow model trails Alignscore, MiniCheck, and the ChatGPT-DA metrics, suggesting that these may be more nuanced. Overall, our results establish that one can achieve performance that is at least generally competitive with SOTA factuality consistency metrics using only information captured by shallow features. [^1] This does not imply that existing metrics are overly or exclusively relying on such cues, only that they might.
@@ 4 What do automatic factuality scores measure?
We next attempt to characterize what factuality metrics measure. First, we examine the correlations between shallow features and the scores produced by these metrics (4.1). Then (4.2), we evaluate the sensitivity of the metrics to controlled changes in summary faithfulness, measuring how effectively they respond to corrections of factual inaccuracies. This allows us to assess their ability to detect and reflect improvements in factual consistency.
4.1 Predicting Automatic Factuality Measures with Shallow Features
Figure 4.1: Spearman correlations between actual and predicted metric scores from an MLP model trained to replicate each metric using shallow features as inputs. Results show moderate-to-strong correlations, indicating potential reliance on superficial text attributes.
We aim to evaluate the extent to which metrics rely on shallow features. To this end we first assess whether we can replicate the scores assigned to summaries by SOTA factuality consistency metrics using shallow predictors only.
To explore this, we again train an MLP model that takes these shallow features as inputs. However, instead of predicting factual inconsistency labels directly (as in the preceding section), we train the model to predict the corresponding scores assigned by the SOTA factuality metrics. In other words, we ask whether such metrics could, in principle, be a function solely of heuristic indicators. Figure 4.1 reports Spearman rank correlations between the predicted scores from the MLP and the metric scores on the test set.
These results show a moderate-to-strong correlation between the predicted scores of the MLP model, trained on shallow features, and the actual metric scores. Note, UniEval and ChatGPT exhibit the weakest correlation among the metrics. These correlations alone do not provide definitive evidence that the metrics necessarily over-rely on shallow indicators; however, they do suggest that such reliance may be a possibility. To try and establish a causal link (i.e., whether such metrics are sensitive to heuristic indicators rather than underlying changes in factual inconsistencies) we next perform a series of controlled manipulations to summaries and measure their responsiveness to these.
4.2 Measuring Metric Sensitivities to Controlled Manipulations
Here we examine the sensitivity of automatic factuality metrics to factual corrections and benign modifications. We aim to clarify whether metrics are responsive to changes in factual consistency, and/or if they are unduly influenced by superficial adjustments with no bearing on factual consistency.
Figure 4.2: Score differences for each metric between the original summary and summaries edited for factual accuracy by humans (shown in green) and other benign edits (shown in blue).
For this analysis, we focus on a subset of summaries manually labelled as inconsistent by human annotators. For each of these, we have a paired, minimally edited version that has been corrected for consistency. These inconsistent/corrected summary pairs were released with the Genaudit dataset (Krishna et al., 2024). Using these, we can measure metric sensitivity to factual consistency (the actual quality of interest).
We are also interested in score variation that owes to superfluous factors. To this end, we prompt GPT-4 to create versions of summaries modified to vary in targeted ways that should be independent of factual consistency. For this we use several prompts requesting different transformations to summaries: Adding least relevant sentences from the source to extend summaries (added source text); Reducing vocabulary diversity (less diverse); Introducing negation words in a meaning-preserving way (negated); Paraphrasing; Shuffling sentence order (shuffled), and; Simplifying text (simplified) or replacing random words in the summary with a synonym (synonym replacement). We provide prompts for summary edits in Appendix C.1.

### v4

@@ Abstract
Modern LLMs can now produce highly readable abstractive summaries, to the point that traditional automated metrics for evaluating summary quality, such as ROUGE, have saturated. However, LLMs still sometimes introduce inaccuracies into summaries, i.e., information inconsistent with or unsupported by the corresponding source. Measuring the occurrence of these often subtle factual inconsistencies automatically has proved challenging. This in turn has motivated development of metrics intended to measure the factual consistency of generated summaries against sources. But are these approaches measuring what they purport to? Or are they mostly exploiting artifacts? In this work, we stress test a range of automatic factuality metrics, including specialized models and LLM-based prompting methods, to probe what they actually capture. Using a shallow classifier to separate “easy” examples for factual evaluation where surface features suffice from “hard” cases requiring deeper reasoning, we find that all metrics show substantial performance drops on the latter. Furthermore, some metrics are more sensitive to benign, fact-preserving edits than to factual corrections. Building on this observation, we demonstrate that most automatic factuality metrics can be gamed, i.e., their scores can be artificially inflated by appending innocuous, content-free sentences to summaries. Among the metrics tested, the LLM-based ChatGPT-DA approach is the most robust and reliable. However, this comes with a notable caveat: Prompting LLMs to assess factuality may overly rely on their parametric knowledge rather than the provided reference when making judgments. Taken together, our findings call into question the reliability of current factuality metrics and prompt a broader reflection on what these metrics are truly measuring. We conclude with concrete recommendations for improving both benchmark design and metric robustness, particularly in light of their vulnerability to superficial manipulations.
@@ 1 Introduction
LLMs are strong abstractive summarizers (goyal2022news; zhang2024benchmarking), but they are not infallible. Even the largest, most capable models sometimes introduce subtle “hallucinations” (or “confabulations”) into summaries that are unsupported by or in contradiction to the corresponding input document (zhang2024benchmarking; tang-etal-2024-tofueval; ramprasad-etal-2024-evaluating). Such behavior is especially problematic in domains such as medicine or law, where inaccurate information could translate into meaningfully negative consequences for individuals.
@@
However, manually evaluating model outputs’ factual consistency with respect to references is expensive, time-consuming, and impractical to scale. This has motivated development of automated methods that score generated summaries for consistency with respect to reference documents. Such metrics have been operationalized using a range of techniques, including entailment (SummaC; laban2022summac), QA models (QuestEval; scialom2021questeval), specialized models explicitly trained to score source-summary pairs (UniEval; zhong2022towards, Alignscore; zha2023alignscore, MiniCheck; tang2024minicheck), and more recently, LLM-based methods that rely on prompting LLMs (ChatGPT-DA; wang2023chatgpt; luo2023chatgpt).
Figure 1: Many methods have been proposed to automatically evaluate the factual consistency of summaries with respect to inputs. In this work we critically evaluate such approaches, e.g., by measuring their sensitivity to various manipulations, as shown here.
Most of these metrics have been evaluated against human benchmark assessments (binary labels or Likert scores) of factual consistency (maynez-etal-2020-faithfulness; fabbri2021summeval; laban2022summac; honovich2022true; wang2022analyzing; gao2023reference; tang2024minicheck).
@@
These assessments have established that automated factuality metrics correlate with human evaluations to varying degrees. But are such approaches actually attuned to the subtle question of factual consistency between inputs and outputs, or are they merely relying on shallow heuristic signals to make their judgments?
If metrics do rely on shallow heuristic patterns, this raises questions about their reliability. It also makes them vulnerable to “gaming”: Summaries or claims can be crafted to exploit heuristic shortcuts and artificially inflate scores without genuinely improving factual accuracy. Alternatively, LLM-based approaches may overly rely on their internal (parametric) knowledge to assess summaries, rather than assessing output factuality with respect to the source. In this work, we stress test a diverse set of SOTA factual consistency metrics on context-claim pairs that yield continuous factual consistency scores. Our analyses offer the following empirical results as our main contributions.
Metrics struggle when factuality requires reasoning beyond shallow cues. In Section 3 we train a shallow MLP on surface-level heuristics and use its prediction confidence to categorize summaries as easy, moderate, or hard to assess for factuality. We find that all automated metrics exhibit substantial performance drops from easy to hard examples, suggesting they may be relying on superficial cues rather than nuanced reasoning.
Factuality metrics are somewhat responsive to factual corrections, but also often sensitive to irrelevant (benign) modifications to summaries. In Section 4.1, we evaluate whether factuality metrics distinguish genuine factual corrections from inconsequential edits using a dataset of annotated summary pairs (krishna2024genaudit) which comprise an inconsistent summary and a minimally edited, faithful revision. Ideally, metrics should assign higher scores to the corrected versions while remaining stable under benign, fact-preserving edits (e.g., paraphrasing). We find that while both specialized and prompt-based metrics respond to factual corrections, many—especially NLI-based SummaC and some specialized models (UniEval and Alignscore)—are overly sensitive to benign edits. In fact, some metrics exhibit greater score shifts from superficial changes than from actual factual improvements. By contrast, ChatGPT-DA consistently ranks faithful revisions higher while remaining robust to benign perturbations, suggesting that it may be a more reliable choice.
@@
(Some) factual consistency metrics are gameable. If factuality metrics rely at least partially on superficial cues, this suggests that we should be able to “game” them by inserting such cues to inflate scores assigned to model outputs. And indeed in Section 5 we find that inserting superfluous, innocuous phrases—either alone or appended to claims——can significantly inflate factuality scores. For NLI-based and specialized metrics, these artificial boosts often exceed the score gains achieved by genuinely more factual models (see Figure 6). ChatGPT-DA, however, shows minimal sensitivity to such manipulations, again suggesting that prompting LLMs for factuality assessment may offer more robust and reliable evaluations. However, this comes with a caveat.
GPT sometimes over relies on internal knowledge when evaluating summaries—even when a reference source is provided. ChatGPT-DA may default to its internal (parametric) knowledge rather than relying on the provided source when assessing factuality consistency (see Section 4.2). This tendency appears especially pronounced when the reference document contradicts the model’s parametric knowledge, indicating that its reliability may be compromised in cases involving myths, rare or updated facts, or information that conflicts with what the model “believes”. This provides a caveat to our other findings that suggest LLM based assessments may be more reliable than bespoke models: LLM-based evaluations may reflect what the (evaluator) model “knows” rather than what the source says, limiting reliability in grounded factuality assessments.
Practical recommendations. Our analyses reveal that while specialized metrics perform well overall, they are vulnerable to benign edits and adversarial manipulation, which may limit their reliability. ChatGPT-DA offers a more robust alternative, but may emphasize consistency with its parametric knowledge rather than input sources, raising concerns about its grounding behavior. We therefore recommend caution when applying such metrics in domains involving myths, misinformation, or uncommon facts. Finally, we highlight the need for benchmarks that capture hallucination severity and for methods that incorporate saliency-aware supervision to improve metric reliability (Section 7)
@@ 2.1 Benchmark Datasets
Automatic factuality metrics for fact verification have been evaluated using various benchmarks, predominantly based on news sources (hermann2015teaching; narayan-etal-2018-dont). The AggreFact dataset (tang2022understanding) consolidates several benchmarks on fine-tuned model-generated summaries from such sources (maynez-etal-2020-faithfulness; wang2020asking; pagnoni2021understanding; fabbri2021summeval; honovich2022true; laban2022summac). Similarly, datasets like FacEval (wang2022analyzing) and ReferenceMatters (gao2023reference) benchmark dialogue summarization of fine-tuned models on sources from SAMSum (gliwa-etal-2019-samsum) and DialogSum (chen-etal-2021-dialogsum).
These datasets primarily focus on summaries from fine-tuned models predating modern LLMs, which excel in zero-shot summarization (goyal2022news). Consequently, factuality metrics evaluated on these benchmarks may not generalize to SOTA LLM outputs, which exhibit different error patterns (tang2022understanding). To address this, tang2024minicheck introduced LLM-AggreFact, a fact-checking dataset from recent LLMs across diverse domains and dataset types. Similarly, krishna2024genaudit released the GenAudit dataset with factuality annotations for LLM summaries in news, Reddit, and clinical settings, while ramprasad-etal-2024-analyzing focus on LLM generated dialogue summaries in the LLM-dialogue dataset.
For our analysis we use all of the above benchmarks to capture a wide range of error types. For fine-tuned model summaries, we use AggreFact for news and FacEval for dialogues. For LLM-generated summaries, we rely on LLM-AggreFact, GenAudit, and LLM-dialogue. We note that each benchmark consolidates multiple datasets and to ensure clean separation of distributions, we avoid overlapping datasets between our test and development splits. Since AggreFact primarily includes older fine-tuned models, we incorporate its dev set into our development data and reserve newer benchmarks for evaluation. Specifically, our dev set includes summaries from the AggreFact dev split, as well as XSUM and CNNDM examples from Genaudit, ensuring no overlap with test data. All remaining datasets are evaluated using their respective test splits. We provide a detailed breakdown of our dev and test splits in Appendix 8.
@@ 2.2 Automatic Factuality Metrics for Summarization
For QA-based metrics we use QuestEval (scialom2021questeval). As an NLI-based metric we use SummaC-Conv (laban2022summac). We use three specialized models for evaluation: UniEval, AlignScore and MiniCheck. UniEval zhong2022towards reframes NLG evaluation as a Boolean QA task and uses T5 (raffel2020exploring) to score different dimensions. AlignScore (zha2023alignscore) evaluates summaries by combining an alignment function—a RoBERTa model (liu2019roberta) fine-tuned on diverse tasks—with a splitting and aggregation strategy. MiniCheck(tang2024minicheck) uses a Flan-T5 model (chung2022scaling) fine-tuned on a synthetic dataset created by the authors.
We also use GPT-4o-mini to score the factual consistency of summaries based on a direct assessment (DA) prompt template from wang2023chatgpt.
@@
QuestEval (scialom2021questeval) QA
SummaC-Conv (laban2022summac) NLI
UniEval (zhong2022towards) specialized model
AlignScore (zha2023alignscore) specialized model
MiniCheck (tang2024minicheck) specialized model
ChatGPT-DA (wang2023chatgpt) Prompt / LLM
@@
3 Do metrics infer “Factuality” from Superficial Features?
The degree to which a claim is faithful to the source text is a subtle question that demands understanding the content in both texts to determine if they are consistent. We would therefore expect that metrics capable of measuring factual consistency in general would need to capitalize on information beyond the superficial features of source and claim texts just discussed.
To investigate the extent to which shallow features explain metric behavior, we train an MLP classifier to predict binary human factuality labels on a development set using only surface-level features. [^1] See Appendix 9 for feature details and Appendix 10 for MLP details. We then apply the trained model to an evaluation set and categorize summaries into three difficulty levels—easy, medium, and hard—based on prediction accuracy and confidence. Confidence is measured as the absolute deviation of the predicted probability from 0.5: Lower values indicate greater uncertainty. We classify a summary as easy if the prediction is correct with high confidence (top 80% of confidence scores), and medium if correct with lower confidence. Incorrect predictions are designated medium if confidence is low (bottom 20%) and hard otherwise. The idea here is that some examples can be readily classified as “factual” (or not) using shallow features like word overlap; these are “easy” examples. By contrast, “medium” and “hard” categories capture examples that are difficult to classify with respect to factuality using these shallow features.
Figure 2: Summaries are categorized as easy, medium, or hard based on prediction accuracy and confidence from a shallow MLP. While specialized metrics perform best on easy examples, their performance declines on hard cases. UniEval, MiniCheck and ChatGPT-DA show greater robustness in more challenging settings
Figure 2 reports the AUC for metrics across summary difficulty levels. All metrics perform best on easy summaries, where shallow features reliably predict human judgments. Performance declines on medium and hard examples, with older metrics like QuestEval and SummaC-Conv showing sharp drops, suggesting reliance on superficial cues. Specialized and prompt-based metrics (e.g., UniEval, AlignScore, MiniCheck, ChatGPT-DA) are more robust, maintaining higher AUC on medium cases. However, even these models struggle on hard summaries, indicating that current metrics, despite being trained for consistency, remain limited in the absence of shallow cues.
@@ 4 What do automatic factuality scores measure?
We next examine the reliability of factuality metrics. In Section 4.1 we evaluate their sensitivity to factual corrections versus benign (fact-preserving) edits. We find that only ChatGPT-DA consistently distinguishes between the two, showing both sensitivity and robustness. In Section 4.2, we evaluate ChatGPT-DA’s reliance on source content by scoring summaries against counterfactual references that conflict with its parametric knowledge. Our results show that consistency evaluation performance degrades when presented with counterfactual input documents, indicating limited grounding in the provided sources.
4.1 Measuring Metric Sensitivities to Controlled Manipulation
We want to determine whether metrics respond specifically to changes in factual consistency or are influenced by superficial edits unrelated to fidelity. We focus on a subset of summaries from the GenAudit dataset (krishna2024genaudit), each manually labeled as inconsistent and paired with a minimally edited, corrected version. These pairs allow us to directly assess whether metrics are sensitive to the consistency improvements that matter most.
Figure 3: Score differences for each metric between the original summary and summaries edited for factual accuracy by humans (shown in green) and other benign edits (shown in blue).
We are also interested in score variation that owes to superfluous factors. To this end, we prompt GPT-4 to generate versions of summaries modified in targeted, fact-preserving ways intended to be independent of factual consistency. We use several prompts that request benign transformations like paraphrasing, simplification, and rewording. [^2] Prompts for benign summary edits are provided in Appendix 11. While these edits are designed to preserve meaning, using GPT may occasionally introduce factual inconsistencies, though we believe such cases are rare.